# TwinFlow: Realizing One-step Generation on Large Models with Self-adversarial Flows

### v1

@@ Abstract
[Uncaptioned image]Models: https://huggingface.co/inclusionAI/TwinFlow
@@ 1 Introduction
Modern generative paradigms—including diffusion (Ho et al., 2020; Song et al., 2020a), flow matching (Lipman et al., 2022; Ma et al., 2024), and consistency models (Song et al., 2023; Lu & Song, 2024)—have achieved state-of-the-art performance, forming the backbone of leading image and video generation systems (Peebles & Xie, 2023; Ho et al., 2022; Chen et al., 2024c; Xie et al., 2024a). Despite their success, these methods share a critical drawback: both training and sampling demand substantial computational resources. This challenge is magnified in the era of large-scale models (ModelTC, 2025). For these systems, efficient sampling is paramount, as the continuous, long-term cost of inference often surpasses the one-time training cost, directly impacting their economic viability and practical deployment (ModelTC, 2025; Xie et al., 2025a).
Numerous research efforts aim to accelerate generative inference by reducing the number of sampling steps. Early single-step methods, like Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), often suffered from unstable training dynamics. To accelerate the multi-step diffusion models, various distillation techniques have been introduced. These range from student-teacher methods, where a compact model learns to emulate a larger one in fewer steps (Salimans & Ho, 2022; Meng et al., 2022), to distribution matching distillation (e.g., DMD variants (Yin et al., 2024b; a)), which uses adversarial training to directly align the model’s output distribution with the real data. In parallel, a powerful new paradigm of consistency models (Song et al., 2023) and their variants, such as LCM (Luo et al., 2023) and PCM (Wang et al., 2024a), has emerged, explicitly designed for high-quality generation in very few steps.
Despite their progress, existing few-step methods face a difficult trade-off between simplicity, efficiency, and quality. (a) Complexity and instability: As summarized in Tab. 1, adversarial methods such as GANs and DMD require auxiliary networks (e.g., discriminators) or frozen teacher models. This not only introduces training instability and sensitivity to hyperparameters but also increases architectural complexity and memory overhead (c.f. Fig. 2(b)), hindering their scalability to large models. (b) Performance degradation: Conversely, methods that train from scratch without adversarial guidance (Luo et al., 2023; Yin et al., 2024a), such as consistency models, often exhibit a sharp decline in quality at very low NFEs (< 4) (Chen et al., 2025b). In summary, we posit that existing methods either suffer from training instability, or require additional/frozen models (see our Tab. 1), which limits their simplicity and scalability in training large models.
@@
Diffusion distillation (Salimans & Ho, 2022) few-step 0 1
@@ 2 Preliminaries
A recent framework, RCGM (Sun & Lin, 2025), introduces a unified formulation for the any-step generation framework, covering paradigms like multi-step generative models (Ho et al., 2020; Song et al., 2020b; Lipman et al., 2022) and few-step generative models (Song et al., 2023; Lu & Song, 2024; Frans et al., 2024; Geng et al., 2025; Sun et al., 2025).
@@
Few-step models are also instances of the RCGM framework, typically corresponding to $N=1$ case: (1) setting $t_{1}=t-\Delta t,\Delta t\to 0$ and $t_{2}=0$ recovers the objective for consistency models (Song et al., 2023; Lu & Song, 2024); (2) setting $t_{1}\in(t_{2},t)$ corresponds to shortcut models (Frans et al., 2024), where the predict function is defined as $\boldsymbol{f}(\mathbf{x}_{t},r)\leftarrow\boldsymbol{f}(\mathbf{x}_{t},s)+\boldsymbol{f}(\mathbf{x}_{r},s)$ and $r=t_{2}\in[0,t]$; (3) setting $t_{1}=t-\Delta t$ (with $\Delta t\to 0$) yields the MeanFlow objective (Geng et al., 2025).
@@ 3.1 Twin Trajectory for Self-adversarial Training
Specifically, we task the network to learn the generative path to its own outputs. We take a fake sample $\mathbf{x}^{\mathrm{fake}}$ generated by the model, i.e. $\mathbf{x}^{\mathrm{fake}}=\hat{\mathbf{x}_{t}}=\boldsymbol{F}_{\theta}(\mathbf{x}_{t}^{\mathrm{real}},t)$, and construct a corresponding “fake trajectory”, in which its perturbed version is defined as $\mathbf{x}^{\mathrm{fake}}_{t^{\prime}}=\alpha(t^{\prime})\mathbf{z}^{\mathrm{fake}}+\gamma(t^{\prime})\mathbf{x}^{\mathrm{fake}}$, $\mathbf{z}^{\mathrm{fake}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and $t^{\prime}\sim\mathrm{U}(0,1)$. Here $\mathbf{z}^{\mathrm{fake}}$ is a different noise, which does not need to be the same as $\mathbf{z}$. The network is then trained with the following flow matching objective on this trajectory, using negative time inputs $t^{\prime}\in[-1,0]$:
@@ 3.2 Rectifying Real Trajectory via Velocity Matching
$=\nabla_{\boldsymbol{\theta}}{\mathbb{E}}_{\mathbf{x}_{t},\mathbf{z},t}\left[\frac{\partial\log p_{\mathrm{fake}}(\mathbf{x}_{t})}{\partial\mathbf{x}_{t}}\;\frac{\partial\mathbf{x}_{t}}{\partial\boldsymbol{\theta}}-\frac{\partial\log p_{\mathrm{real}}(\mathbf{x}_{t})}{\partial\mathbf{x}_{t}}\;\frac{\partial\mathbf{x}_{t}}{\partial\boldsymbol{\theta}}\right]$
$=\nabla_{\boldsymbol{\theta}}{\mathbb{E}}_{\mathbf{x}_{t},\mathbf{z},t}\left[\Big(\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{\mathrm{fake}}(\mathbf{x}_{t})-\nabla_{\mathbf{x}_{t}}\log p_{\mathrm{real}}(\mathbf{x}_{t})}_{\displaystyle\mathbf{s}_{\mathrm{fake}}(\mathbf{x}_{t})-\mathbf{s}_{\mathrm{real}}(\mathbf{x}_{t})}\Big)\frac{\partial\mathbf{x}_{t}}{\partial\boldsymbol{\theta}}\right]\,$, (4)
@@
$\textstyle\nabla_{\boldsymbol{\theta}}D_{\mathrm{KL}}$ $=\nabla_{\boldsymbol{\theta}}{\mathbb{E}}_{\mathbf{x}_{t},\mathbf{z},t}\left[\left(-\frac{\mathbf{x}_{t}+(1-t)\cdot\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-t})}{t}-(-\frac{\mathbf{x}_{t}+(1-t)\cdot\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}t})}{t})\right)\frac{\partial\mathbf{x}_{t}}{\partial\boldsymbol{\theta}}\right]$
$=\nabla_{\boldsymbol{\theta}}{\mathbb{E}}_{\mathbf{x}_{t},\mathbf{z},t}\left[-\frac{(1-t)}{t}\cdot\Big(\underbrace{\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-t})}_{\displaystyle\mathbf{v}_{\mathrm{fake}}(\mathbf{x}_{t},-t)}-\underbrace{\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}t})}_{\displaystyle\mathbf{v}_{\mathrm{real}}(\mathbf{x}_{t},t)}\Big)\frac{\partial\mathbf{x}_{t}}{\partial\boldsymbol{\theta}}\right]\,$, (6)
@@
$\Delta_{\mathbf{v}}(\mathbf{x}_{t}):=\mathbf{v}_{\mathrm{fake}}(\mathbf{x}_{t},-t)-\mathbf{v}_{\mathrm{real}}(\mathbf{x}_{t},t)\,.$ (7)
@@
To derive the rectification loss, we first instantiate the gradient (6) using the setup in Sec. 3.1. In this setting, the network’s prediction $\hat{\mathbf{x}}_{t}$ serves as the clean example, and consequently, the perturbed variable $\mathbf{x}_{t}$ in (6) corresponds to the fake sample $\mathbf{x}^{\mathrm{fake}}_{t^{\prime}}$. The velocity difference $\Delta_{\mathbf{v}}(\mathbf{x}_{t})$ defined in (7) is therefore instantiated as $\Delta_{\mathbf{v}}(\mathbf{x}^{\mathrm{fake}}_{t^{\prime}})=\mathbf{v}_{\mathrm{fake}}(\mathbf{x}^{\mathrm{fake}}_{t^{\prime}},-t^{\prime})-\mathbf{v}_{\mathrm{real}}(\mathbf{x}^{\mathrm{fake}}_{t^{\prime}},t^{\prime})$.
@@
$\frac{\partial\mathbf{x}^{\mathrm{fake}}_{t^{\prime}}}{\partial\boldsymbol{\theta}}=\frac{\partial(\alpha(t^{\prime})\mathbf{z}^{\mathrm{fake}}+\gamma(t^{\prime})\mathbf{x}^{\mathrm{fake}})}{\partial\boldsymbol{\theta}}=\frac{\partial(\alpha(t^{\prime})\mathbf{z}^{\mathrm{fake}}+\gamma(t^{\prime})\hat{\mathbf{x}_{t}})}{\partial\boldsymbol{\theta}}=\frac{\partial\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t}^{\mathrm{real}},t)}{\partial\boldsymbol{\theta}}\,.$ (8)
@@
$\mathcal{L}(\boldsymbol{\theta})_{\mathrm{rectify}}={\mathbb{E}}_{\mathbf{x}_{t},\mathbf{x}^{\mathrm{fake}},\mathbf{z}^{\mathrm{fake}},t^{\prime}}\left[d\left(\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t}^{\mathrm{real}},t),\mathrm{sg}\left[\Delta_{\mathbf{v}}(\mathbf{x}^{\mathrm{fake}}_{t^{\prime}})+\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t}^{\mathrm{real}},t)\right]\right)\right]\,$, (9)
@@ 4.1 Experimental Setup
  • Image generation on multimodal generative models. We conduct evaluations on unified multi-modal models (i.e. takes both texts and images as conditions and capable of generating texts and images). (1) Network architectures: We conducted LoRA (Hu et al., 2022) (Tab. 2) and full-parameter training (Tab. 3) of TwinFlow on Qwen-Image. We also do full-parameter training experiments on OpenUni-512 (Wu et al., 2025c). (2) Benchmarks: Following recent works (Pan et al., 2025; Chen et al., 2025a; Deng et al., 2025; Wu et al., 2025a), we use benchmarks in text-to-image generation tasks. For text-to-image generation, we use GenEval, DPG-Bench (Hu et al., 2024), and WISE (Niu et al., 2025). Other training settings are detailed in App. B.1.
@@ 4.2 Image Generation on Multimodal Generative Models
As shown in Tab. 2, Qwen-Image-TwinFlow achieves a score of 0.86 on GenEval and 86.52% on DPG-Bench with just 1-NFE, closely matching the original model’s performance at 100-NFE. Compared to Qwen-Image-Lightning (ModelTC, 2025), a 4-step distilled model, our model surpasses it on GenEval and WISE with only 1-NFE. Furthermore, our model outperforms Qwen-Image-RCGM (Sun & Lin, 2025) on both GenEval and DPG-Bench under 1-NFE and 2-NFE settings, with notable improvements of 0.34 on GenEval, 27.0% on DPG-Bench, and 0.25 on WISE under the 1-NFE setting.
We also benchmark Qwen-Image-TwinFlow against other prominent multi-step unified multimodal generative models, such as MetaQuery-XL (Pan et al., 2025), BLIP3-o-8B (Chen et al., 2025a), and Bagel (Deng et al., 2025). Our model consistently surpasses these baselines with 1 or 2-NFE across all evaluation metrics. Beyond Qwen-Image, we also apply TwinFlow to OpenUni (Wu et al., 2025c), achieving GenEval of 0.80 and DPG-Bench of 76.40 under the 1-NFE setting, which is also close to its original performance. These findings underscore the versatility and effectiveness of TwinFlow across different architectures and scales.
@@
OpenUni-RCGM-512 (Sun & Lin, 2025) 2 0.85 80.15 0.50
OpenUni-RCGM-512 (Sun & Lin, 2025) 1 0.80 76.40 0.45
@@
Qwen-Image-RCGM (Sun & Lin, 2025) 2 0.82 84.09 0.50
Qwen-Image-RCGM (Sun & Lin, 2025) 1 0.52 59.50 0.30
@@
    Performance superiority: With this unified design, TwinFlow outperforms all baselines on Qwen-Image-20B. Notably, it achieves superior generation quality with just 1-2 NFE compared to sCM (Lu & Song, 2024) and MeanFlow (Geng et al., 2025) operating at 8 NFE.

### v2

@@ Abstract
[Uncaptioned image]Models: https://huggingface.co/collections/inclusionAI/twinflow
@@ 1 Introduction
Modern generative paradigms—including diffusion (Ho et al., 2020; Song et al., 2020a), flow matching (Lipman et al., 2022; Ma et al., 2024), and consistency models (Song et al., 2023; Lu and Song, 2024)—have achieved state-of-the-art performance, forming the backbone of leading image and video generation systems (Peebles and Xie, 2023; Ho et al., 2022; Chen et al., 2024c; Xie et al., 2024a). Despite their success, these methods share a critical drawback: both training and sampling demand substantial computational resources. This challenge is magnified in the era of large-scale models (ModelTC, 2025). For these systems, efficient sampling is paramount, as the continuous, long-term cost of inference often surpasses the one-time training cost, directly impacting their economic viability and practical deployment (ModelTC, 2025; Xie et al., 2025a).
Numerous research efforts aim to accelerate generative inference by reducing the number of sampling steps. Early single-step methods, like Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), often suffered from unstable training dynamics. To accelerate the multi-step diffusion models, various distillation techniques have been introduced. These range from student-teacher methods, where a compact model learns to emulate a larger one in fewer steps (Salimans and Ho, 2022; Meng et al., 2022), to distribution matching distillation (e.g., DMD variants (Yin et al., 2024b; a)), which uses adversarial training to directly align the model’s output distribution with the real data. In parallel, a powerful new paradigm of consistency models (Song et al., 2023) and their variants, such as LCM (Luo et al., 2023) and PCM (Wang et al., 2024a), has emerged, explicitly designed for high-quality generation in very few steps.
Despite their progress, existing few-step methods face a difficult trade-off between simplicity, efficiency, and quality. (a) Complexity and instability: As summarized in Tab. 1, adversarial methods such as GANs and DMD require auxiliary networks (e.g., discriminators) or frozen teacher models. This not only introduces training instability and sensitivity to hyperparameters but also increases architectural complexity and memory overhead (c.f. Fig. 2(b)), hindering their scalability to large models. (b) Performance degradation: Conversely, methods that train from scratch without adversarial guidance (Luo et al., 2023; Yin et al., 2024a), such as consistency models, often exhibit a sharp decline in quality at very low NFEs (< 4) (Chen et al., 2025c). In summary, we posit that existing methods either suffer from training instability, or require additional/frozen models (see our Tab. 1), which limits their simplicity and scalability in training large models.
@@
Diffusion distillation (Salimans and Ho, 2022) few-step 0 1
@@ 2 Preliminaries
A recent framework, RCGM (Sun and Lin, 2025), introduces a unified formulation for the any-step generation framework, covering paradigms like multi-step generative models (Ho et al., 2020; Song et al., 2020b; Lipman et al., 2022) and few-step generative models (Song et al., 2023; Lu and Song, 2024; Frans et al., 2024; Geng et al., 2025; Sun et al., 2025).
@@
Few-step models are also instances of the RCGM framework, typically corresponding to $N=1$ case: (1) setting $t_{1}=t-\Delta t,\Delta t\to 0$ and $t_{2}=0$ recovers the objective for consistency models (Song et al., 2023; Lu and Song, 2024); (2) setting $t_{1}\in(t_{2},t)$ corresponds to shortcut models (Frans et al., 2024), where the predict function is defined as $\boldsymbol{f}(\mathbf{x}_{t},r)\leftarrow\boldsymbol{f}(\mathbf{x}_{t},s)+\boldsymbol{f}(\mathbf{x}_{r},s)$ and $r=t_{2}\in[0,t]$; (3) setting $t_{1}=t-\Delta t$ (with $\Delta t\to 0$) yields the MeanFlow objective (Geng et al., 2025).
@@ 3.1 Twin Trajectory for Self-adversarial Training
Specifically, we task the network to learn the generative path to its own outputs. We take a fake sample $\mathbf{x}^{\mathrm{fake}}$ generated by the model, i.e. $\mathbf{x}^{\mathrm{fake}}=\hat{\mathbf{x}_{t}}=\mathbf{x}_{t}^{\mathrm{real}}-\boldsymbol{F}_{\theta}(\mathbf{x}_{t}^{\mathrm{real}},r)\cdot t$ ($t=1,r=0$ are used in implementation, i.e. $\mathbf{x}^{\mathrm{fake}}=\mathbf{z}-\boldsymbol{F}_{\theta}(\mathbf{z},0)$), and construct a corresponding “fake trajectory”, in which its perturbed version is defined as $\mathbf{x}^{\mathrm{fake}}_{t^{\prime}}=\alpha(t^{\prime})\mathbf{z}^{\mathrm{fake}}+\gamma(t^{\prime})\mathbf{x}^{\mathrm{fake}}$, $\mathbf{z}^{\mathrm{fake}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and $t^{\prime}\sim\mathrm{U}(0,1)$. Here $\mathbf{z}^{\mathrm{fake}}$ is a different noise, which does not need to be the same as $\mathbf{z}$. The network is then trained with the following flow matching objective on this trajectory, using negative time inputs $t^{\prime}\in[-1,0]$:
@@ 3.2 Rectifying Real Trajectory via Velocity Matching
$={\mathbb{E}}_{\mathbf{x}_{t},\mathbf{z},t}\left[\frac{\partial\log p_{\mathrm{fake}}(\mathbf{x}_{t})}{\partial\mathbf{x}_{t}}\;\frac{\partial\mathbf{x}_{t}}{\partial\boldsymbol{\theta}}-\frac{\partial\log p_{\mathrm{real}}(\mathbf{x}_{t})}{\partial\mathbf{x}_{t}}\;\frac{\partial\mathbf{x}_{t}}{\partial\boldsymbol{\theta}}\right]$
$={\mathbb{E}}_{\mathbf{x}_{t},\mathbf{z},t}\left[\Big(\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{\mathrm{fake}}(\mathbf{x}_{t})-\nabla_{\mathbf{x}_{t}}\log p_{\mathrm{real}}(\mathbf{x}_{t})}_{\displaystyle\mathbf{s}_{\mathrm{fake}}(\mathbf{x}_{t})-\mathbf{s}_{\mathrm{real}}(\mathbf{x}_{t})}\Big)\frac{\partial\mathbf{x}_{t}}{\partial\boldsymbol{\theta}}\right]\,$, (4)
@@
$\textstyle\nabla_{\boldsymbol{\theta}}D_{\mathrm{KL}}$ $={\mathbb{E}}_{\mathbf{x}_{t},\mathbf{z},t}\left[\left(-\frac{\mathbf{x}_{t}+(1-t)\cdot\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-t})}{t}-(-\frac{\mathbf{x}_{t}+(1-t)\cdot\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}t})}{t})\right)\frac{\partial\mathbf{x}_{t}}{\partial\boldsymbol{\theta}}\right]$
$={\mathbb{E}}_{\mathbf{x}_{t},\mathbf{z},t}\left[-\frac{(1-t)}{t}\cdot\Big(\underbrace{\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-t})}_{\displaystyle\mathbf{v}_{\mathrm{fake}}(\mathbf{x}_{t},-t)}-\underbrace{\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}t})}_{\displaystyle\mathbf{v}_{\mathrm{real}}(\mathbf{x}_{t},t)}\Big)\frac{\partial\mathbf{x}_{t}}{\partial\boldsymbol{\theta}}\right]\,$, (6)
@@
$\Delta_{\mathbf{v}}(\mathbf{x}_{t}):=\mathbf{v}_{\mathrm{real}}(\mathbf{x}_{t},t)-\mathbf{v}_{\mathrm{fake}}(\mathbf{x}_{t},-t)\,.$ (7)
@@
To derive the rectification loss, we first instantiate the gradient (6) using the setup in Sec. 3.1. In this setting, the network’s prediction $\hat{\mathbf{x}}_{t}$ serves as the clean example, and consequently, the perturbed variable $\mathbf{x}_{t}$ in (6) corresponds to the fake sample $\mathbf{x}^{\mathrm{fake}}_{t^{\prime}}$. The velocity difference $\Delta_{\mathbf{v}}(\mathbf{x}_{t})$ defined in (7) is therefore instantiated as $\Delta_{\mathbf{v}}(\mathbf{x}^{\mathrm{fake}}_{t^{\prime}})=\mathbf{v}_{\mathrm{real}}(\mathbf{x}^{\mathrm{fake}}_{t^{\prime}},t^{\prime})-\mathbf{v}_{\mathrm{fake}}(\mathbf{x}^{\mathrm{fake}}_{t^{\prime}},-t^{\prime})$.
@@
$\frac{\partial\mathbf{x}^{\mathrm{fake}}_{t^{\prime}}}{\partial\boldsymbol{\theta}}=\frac{\partial(\alpha(t^{\prime})\mathbf{z}^{\mathrm{fake}}+\gamma(t^{\prime})\mathbf{x}^{\mathrm{fake}})}{\partial\boldsymbol{\theta}}=\frac{\partial(\alpha(t^{\prime})\mathbf{z}^{\mathrm{fake}}+\gamma(t^{\prime})\hat{\mathbf{x}_{t}})}{\partial\boldsymbol{\theta}}\propto-\frac{\partial\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t}^{\mathrm{real}},r)}{\partial\boldsymbol{\theta}}$ $\stackrel{{\scriptstyle t=1,\,r=0}}{{=}}-\frac{\partial\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{z},0)}{\partial\boldsymbol{\theta}}\,.$ (8)
@@
$\mathcal{L}(\boldsymbol{\theta})_{\mathrm{rectify}}={\mathbb{E}}_{\mathbf{x}_{t},\mathbf{x}^{\mathrm{fake}},\mathbf{z}^{\mathrm{fake}},t^{\prime}}\left[d\left(\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{z},0),\mathrm{sg}\left[\Delta_{\mathbf{v}}(\mathbf{x}^{\mathrm{fake}}_{t^{\prime}})+\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{z},0)\right]\right)\right]\,$, (9)
@@ 4.1 Experimental Setup
  • Image generation on multimodal generative models. We conduct evaluations on unified multi-modal models (i.e. takes both texts and images as conditions and capable of generating texts and images). (1) Network architectures: We conducted LoRA (Hu et al., 2022) (Tab. 2) and full-parameter training (Tab. 3) of TwinFlow on Qwen-Image. We also do full-parameter training experiments on OpenUni-512 (Wu et al., 2025c). (2) Benchmarks: Following recent works (Pan et al., 2025; Chen et al., 2025b; Deng et al., 2025; Wu et al., 2025a), we use benchmarks in text-to-image generation tasks. For text-to-image generation, we use GenEval, DPG-Bench (Hu et al., 2024), and WISE (Niu et al., 2025). Other training settings are detailed in App. B.1.
@@ 4.2 Image Generation on Multimodal Generative Models
As shown in Tab. 2, Qwen-Image-TwinFlow achieves a score of 0.86 on GenEval and 86.52% on DPG-Bench with just 1-NFE, closely matching the original model’s performance at 100-NFE. Compared to Qwen-Image-Lightning (ModelTC, 2025), a 4-step distilled model, our model surpasses it on GenEval and WISE with only 1-NFE. Furthermore, our model outperforms Qwen-Image-RCGM (Sun and Lin, 2025) on both GenEval and DPG-Bench under 1-NFE and 2-NFE settings, with notable improvements of 0.34 on GenEval, 27.0% on DPG-Bench, and 0.25 on WISE under the 1-NFE setting.
We also benchmark Qwen-Image-TwinFlow against other prominent multi-step unified multimodal generative models, such as MetaQuery-XL (Pan et al., 2025), BLIP3-o-8B (Chen et al., 2025b), and Bagel (Deng et al., 2025). Our model consistently surpasses these baselines with 1 or 2-NFE across all evaluation metrics. Beyond Qwen-Image, we also apply TwinFlow to OpenUni (Wu et al., 2025c), achieving GenEval of 0.80 and DPG-Bench of 76.40 under the 1-NFE setting, which is also close to its original performance. These findings underscore the versatility and effectiveness of TwinFlow across different architectures and scales.
@@
OpenUni-RCGM-512 (Sun and Lin, 2025) 2 0.85 80.15 0.50
OpenUni-RCGM-512 (Sun and Lin, 2025) 1 0.80 76.40 0.45
@@
Qwen-Image-RCGM (Sun and Lin, 2025) 2 0.82 84.09 0.50
Qwen-Image-RCGM (Sun and Lin, 2025) 1 0.52 59.50 0.30
@@
    Performance superiority: With this unified design, TwinFlow outperforms all baselines on Qwen-Image-20B. Notably, it achieves superior generation quality with just 1-2 NFE compared to sCM (Lu and Song, 2024) and MeanFlow (Geng et al., 2025) operating at 8 NFE.